passes through (-4,-2) parrallell to -7x-5y=58
Answer by mukhopadhyay(490)

Line parallel to -7x-5y=58 will have the same slope as this.
Equation of any line parallel to above line is y=-7/5x+k;
The above line passes thru (-4,-2)
=> -2 = (-7/5)(-4) + k
=> k = -2-(28/5)=-38/5
Equation is: y = (-7/5)x -38/5
Same as 5y+7x = -38